A Level Mathematics 9709 Paper 53, May/June 2025

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/53 · May/June 2025

Download PDF
1

For a set of 40 values of x, it is found that Σ(x−k)=836.0,Σ(x−k)2=25410.8, where k is a constant.

a

Given that the mean of these 40 values is 124.0, find the value of k.

[2]
b

Find the standard deviation of these 40 values of x.

[2]
2

At a large college, all students who study Science also study exactly one of Art or Drama or Music. 20% of these students study Art, 45% study Drama and 35% study Music.

a

3 students are selected at random from the students who study Science. Find the probability that at least 1 of these students studies Drama.

[2]
b

10 students are selected at random from the students who study Science. Find the probability that more than 7 study Art or Music.

[3]
3

A fair six-sided dice with faces labelled 1, 2, 3, 4, 5, 6 is thrown repeatedly until a 3 is obtained. The number of throws taken is denoted by the random variable X.

a

Find P(X=8).

[1]
b

Find P(X<9).

[2]
c

Find the probability that a 3 is obtained for the second time before the 6th throw.

[3]
4

84 people attempt a particular puzzle. The times taken, in minutes, to complete the puzzle are recorded. These times are represented in the cumulative frequency graph below.

Figure 1 of 1 for question 4
a

Use the graph to estimate how many people took between 4 and 7.5 minutes to complete the puzzle.

[1]
b

On the grid below, draw a box-and-whisker plot to summarise the information in the cumulative frequency graph.

[4]
5

Bag A contains 6 red marbles, 5 blue marbles and 1 green marble. Bag B contains 5 red marbles and 3 blue marbles. A marble is chosen at random from bag A and placed in bag B. A marble is now chosen at random from bag B.

a

Draw a tree diagram to represent this information, giving the probability on each branch.

[3]
b

Find the probability that both marbles chosen are the same colour.

[2]
c

Find the probability that the marble chosen from bag A is blue, given that the marble chosen from bag B is blue.

[3]
6

A company sells bags of pasta. The masses of large bags of pasta are normally distributed with mean 2.50 kg and standard deviation 0.12 kg.

a

Find the probability that the mass of pasta in a randomly chosen large bag is less than 2.65 kg.

[2]
b

A restaurant manager buys 160 of these large bags of pasta. Find the number of bags for which you would expect the mass of pasta to be more than 1.65 standard deviations above the mean.

[3]
c

The masses of small bags of pasta sold by the company are normally distributed with mean μ kg and standard deviation σ kg. Tests show that 77% of these bags have masses greater than 1.26 kg, and 44% have masses less than 1.35 kg. Find, in either order, the value of μ and the value of σ.

[5]
7

A set of friends consists of 7 men and 4 women. Three of the men are brothers: Ali, Ben and Charlie.

a

Find the number of different arrangements of the 7 men in a line in which Ali and Ben do not stand next to each other.

[3]
b

Find the number of different arrangements of the 7 men and 4 women in a line in which all the men stand together and all the women stand together.

[3]
c

In how many ways can the 7 men and 4 women be divided into a group of 6, a group of 3 and a group of 2 if there are no restrictions?

[2]
d

The 7 men and 4 women are divided at random into a group of 6, a group of 3 and a group of 2. Find the probability that Ali, Ben and Charlie are all in the same group.

[4]

Mark scheme

9709/53 · May/June 2025

Download PDF

1a · 2 marks

  • B1

    4960 – 40k = 836 or 496040−k=83640

  • B1

    k=103.1

1b · 2 marks

  • M1

    [sd2=]25410.840−(83640)2 or [sd2=]622978.440−1242

  • A1

    [sd2=198.46,] sd = 14.1

2a · 2 marks

  • M1

    1−(0.55)3 or 3C1(0.45)(0.55)2+3C2(0.45)2(0.55)+(0.45)3[0.55]0

  • A1

    =0.834,66698000

2b · 3 marks

  • M1

    One term of the form 10Cx(p)x(1−p)10−x,0<p<1,x≠0 or 10.

  • A1

    Correct un-simplified expression.

  • B1

    =0.0996

3a · 1 mark

  • B1

    [P(X=8)=](16)(56)7=0.0465,781251679616

3b · 2 marks

  • M1

    1−(56)d,d=8,9 or full addition of terms

  • A1

    =0.767

3c · 3 marks

  • B1

    Two identified scenarios correct, accept un-simplified calculations (or 1 – 1 identified scenario correct in Method 2).

  • M1

    Four correctly identified scenarios added (or 1 – 2 correctly identified scenarios in Method 2). Condone 2C1=2, 5C1=5 etc.

  • A1

    0.196,7633888

4a · 1 mark

  • B1

    Accept 43 or 44

4b · 4 marks

  • B1

    Median = 4.8 plotted on box.

  • B1

    LQ = 3.7 or 3.8, UQ = 6.5 or 6.6 plotted on box.

  • B1

    Correct end points for whiskers and whiskers drawn

  • B1

    Linear scale and labelled Time in minutes

5a · 3 marks

  • B1

    Correct structure and probabilities for Bag A branches which must be the first set of branches, including labels R, B, G.

  • B1

    Correct structure and probabilities for one Bag B branch, including labels R, B, [G]. Condone inclusion of branches with 0 probability.

  • B1

    Correct structure and probabilities for remaining Bag B branches, including labels R, B, [G]. Condone inclusion of branches with 0 probability.

5b · 2 marks

  • M1

    [P(RR)+P(BB)+P(GG)=]612×69+512×49+112×19

  • A1

    =57108,1936

5c · 3 marks

  • M1

    512×49 or 20108 or 527 or their P(BB) from Q5(b) seen as a numerator of a single fraction. FT tree diagram with values from only 2 bags.

  • B1

    612×39+512×49+112×39 seen as the denominator of a single fraction. 527 or their P(BB) may be seen from Q5b in place of 512×49. FT tree diagram with values from only 2 bags.

  • A1

    =2041

6a · 2 marks

  • M1

    Substituting 2.65, 2.50 and 0.12 into the ±standardisation formula allow σ2 or σ. Condone continuity correction ±0.005.

  • A1

    =0.894

6b · 3 marks

  • M1

    Seeing 2.50+1.65×0.12 or 2.698 and 2.50 and 0.12 substituted into the ±standardisation formula with a direct attempt to find the probability area.

  • A1

    =1−0.9505=0.0495

  • B1

    Expected number=0.0495×160=7.92, so 7 or 8

6c · 5 marks

  • B1

    0.738<z1<0.740 or −0.740<z1<−0.738 seen.

  • B1

    0.150<z2<0.152 or −0.152<z2<−0.150 seen.

  • M1

    Use of the ±standardisation formula once with μ, σ equating to a z-value (not 0.77, 0.23, 0.44, 0.56, 0.7794, 0.5910, 0.6700, 0.7123, 0.2206, 0.4090, 0.3300, 0.2877). No continuity correction, not σ2, not σ.

  • M1

    Solve 2 equations in μ and σ using the elimination method, substitution method or other appropriate approach to obtain values for both μ and σ.

  • A1

    μ=1.37,σ=0.153

7a · 3 marks

  • M1

    7!−k,k an integer ⩾1 (or 5!×p where p an integer >1).

  • M1

    m−6!×n:n=1,2 where m an integer >6! (n=1 can be implied) (or q×6×5,q×6C2,q×6P2,q×6×52, where q an integer >1).

  • A1

    =3600

7b · 3 marks

  • M1

    7!×r, where r is an integer, r>1 or 4!×s, where s is an integer, s>1.

  • M1

    7!×4!×t, where t is an integer, t⩾1, t=1 can be implied.

  • A1

    241920

7c · 2 marks

  • M1

    11Cv×11−vCu×[11−v−uCw], u,v,w=6,3,2,u≠v≠w.

  • A1

    =4620

7d · 4 marks

  • B1

    group of 6: 8C3×5C3×(2C2)=560 (560 seen, accept un-simplified; or 433,0.1212… seen)

  • B1

    group of 3: 8C6(×2C2×3C3)=28 (28 seen accept un-simplified; or 1165,0.00606(06… ) seen)

  • M1

    Probability they are in same group=their(560+28)their (c) or their(560+28)4620 (or their 433+their 1165)

  • A1

    =5884620,755,0.127

Questions

  1. 1
    Question 14 marks
    Mark scheme ›
  2. 2
    Question 25 marks
    Mark scheme ›
  3. 3
    Question 36 marks
    Mark scheme ›
  4. 4
    Question 45 marks
    Mark scheme ›
  5. 5
    Question 58 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›
  7. 7
    Question 712 marks
    Mark scheme ›

More 9709 papers

Other papers from May/June 2025

Solve it on ExamPalStart